A line passes through the point (-2,-5) and is perpendicular to the line -2x+y=9
what is the equation of the line?

Answer by ewatrrr(24785)

 
Hi,
Line perpendicular to the line:
-2x+y=9 OR y = 2x+ 9 slope m = 2
New Line: y = mx + b
y = (-1/2)x + b |Has a slope that is the negative reciprocal of m = 2
y = (-1/2)x + b |Using Pt(-2,-5) to solve for b
-5 = (-1/2)*-2 + b
-6 = b
Equation of the new Line is y = (-1/2)x - 6